In a conventional stack sampling train for particulate matter, water vapor is removed from the sampled gases by condensation and desiccation to protect the gas volume meter. If 90 g of water was removed in a particular test, estimate the partial volume of this water vapor [m3] at the stack conditions of 180°C and 95 kPa.
Answer #1

Moles of water = n = mass/molar mass = 90/18 = 5 moles

Let volume is V liters

Temperature (T) = 180°C = 273 + 180 = 453 K

Pressure (P) = 95 kPa = 95 * 1000 * 9.869 * 10^-6 = 0.938 atm

Gas constant (R) = 0.0821 L atm / mol K

Now, using ideal gas equation,

PV = n R T

V = nRT / P = 5 * 0.0821 * 453 / 0.938 = 198.34 liters

We know that,

1 liter = 0.001 m^3

So, 198.34 L = 198.34 * 0.001 = 0.19834 meter^3 .... Answer

Let me know in case of any question.